OpenStudy (anonymous):

Help! Find the exact value of the expression w/o a calculator? (1/(sec^2 40 degrees)) + (1/csc^2 40 degrees) Can you explain too?

OpenStudy (anonymous):

easy

OpenStudy (anonymous):

lolno

OpenStudy (anonymous):

= cos^24 + si^2 40 = 1

OpenStudy (anonymous):

use x instead of 40 might make it easier

OpenStudy (amistre64):

good thing openstudy isnt a calculator :)

OpenStudy (amistre64):

find 2 degrees that you already know to calculate with

OpenStudy (anonymous):

I was going to say that

OpenStudy (anonymous):

yes - this is a standard trig identity: cos^2x + sin^2x = 1

OpenStudy (anonymous):

I was also going to say find csc(40) * csc (40)

OpenStudy (anonymous):

\[\frac{1}{\sec^2(x)}+\frac{1}{\csc^2(x)}=\cos^2(x)+\sin^2(x)\]

OpenStudy (anonymous):

finally someone saw it lol, nice job satellite :)

OpenStudy (anonymous):

the "40" degrees is a red herring. makes no difference what the input is, you will get 1 in any case

OpenStudy (anonymous):

Tricksters. ty!

OpenStudy (anonymous):

yw from all of us i am sure

OpenStudy (anonymous):

There's more than one way to go about this, but knowing the identity helps

OpenStudy (anonymous):

my scanner isnt as fast as satellite lol :( i need a new scanner >.>

OpenStudy (anonymous):

wow that was fast. mine takes like ten minutes

OpenStudy (anonymous):

not fast enough! haha